# Non-stabilizerness in U(1) lattice gauge theory

### v1

@@ Acknowledgements.
We thank M. Collura, T. Haug and P. Hauke for discussions, and A. Paviglianiti for comments on the manuscript. P.R.N.F thanks A.S. Aramthottil and Konrad Pawlik for the valuable discussion on the numerics. Discussions with B. Damski are also appreciated (J.Z.). E.T. acknowledges J. Mildenberger, G. Chandra Santra, and R. Costa de Almeida for insightful discussions. We acknowledge Polish high-performance computing infrastructure PLGrid for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through PLL/2023/04/016502. The work of P.R.N.F. and that of J.Z. was funded by the National Science Centre, Poland, project 2021/03/Y/ST2/00186 within the QuantERA II Programme (DYNAMITE) that has received funding from the European Union Horizon 2020 research and innovation programme under Grant agreement No 101017733. The study was also partially funded by the ”Research Support Module” as part of the ”Excellence Initiative – Research University” program at the Jagiellonian University in Kraków. M. D.and E. T. were partly supported by the QUANTERA DYNAMITE PCI2022-132919. M. D. was also supported by the EU-Flagship programme Pasquans2, by the PNRR MUR project PE0000023-NQSTI, the PRIN programme (project CoQuS), and the ERC Consolidator grant WaveNets. P. S. T. acknowledges support from the Simons Foundation through Award 284558FY19 to the ICTP. Views and opinions expressed in this work are, however, those of the authors only and do not necessarily reflect those of the European Union, European Climate, Infrastructure and Environment Executive Agency (CINEA), nor any other granting authority Note added: While finalizing this work, we became aware of a study of stabilizer Renyi entropies in the context of PXP models [78]. The two works utilize very different techniques to compute magic and are complementary in terms of parameter regimes. In the only overlapping computation (discussed in Ref. [70]), the agreement between out analytics and the simulations in Ref. [78] is perfect.

### v2

@@ Abstract
Acknowledgments. -We thank M. Collura, T. Haug and P. Hauke for discussions, and A. Paviglianiti for comments on the manuscript. P.R.N.F thanks A.S. Aramthottil and Konrad Pawlik for the valuable discussion on the numerics. Discussions with B. Damski are also appreciated (J.Z.). E.T. acknowledges J. Mildenberger, G. Chandra Santra, and R. Costa de Almeida for insightful discussions. We acknowledge Polish high-performance computing infrastructure PLGrid for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through PLL/2023/04/016502. The work of P.R.N.F. and that of J.Z. was funded by the National Science Centre, Poland, project 2021/03/Y/ST2/00186 within the QuantERA II Programme (DYNAMITE) that has received funding from the European Union Horizon 2020 research and innovation programme under Grant agreement No 101017733. The study was also partially funded by the ”Research Support Module” as part of the ”Excellence Initiative – Research University” program at the Jagiellonian University in Kraków. M. D.and E. T. were partly supported by the QUANTERA DYNAMITE PCI2022-132919. M. D. was also supported by the EU-Flagship programme Pasquans2, by the PNRR MUR project PE0000023-NQSTI, the PRIN programme (project CoQuS), and the ERC Consolidator grant WaveNets. P. S. T. acknowledges support from the Simons Foundation through Award 284558FY19 to the ICTP. Views and opinions expressed in this work are, however, those of the authors only and do not necessarily reflect those of the European Union, European Climate, Infrastructure and Environment Executive Agency (CINEA), nor any other granting authority
Note added: While finalizing this work, we became aware of a study of stabilizer Renyi entropies in the context of PXP models [87]. The two works utilize very different techniques to compute SRE-2 density $m_{2}$ and are complementary in terms of parameter regimes. In the only overlapping computation (discussed in Ref. [70]), the agreement between out analytics and the simulations in Ref. [87] is perfect.
Supplementary material: Nonstabilizerness in U(1) lattice gauge theory
.1 Nonstabilizerness of “rainbow”-like state
Here, we show that our results can also be used to compute the SRE of the “rainbow”-like state
|$\ket{E}=\frac{1}{\sqrt{\chi}}\sum_{f\in F_{L}}(-1)^{|f|}\ket{f}\otimes\ket{f},$ (S.1)
which is an exact volume-law entangled eigenstate at energy $E=0$ in the PXP chain with PBCs and size $N=2L$. Here, $|f|$ denotes the parity of a given configuration. The normalization is given by $\chi=Z_{L}(\beta=0)$. We first notice that applying the operator $\otimes_{i=1}^{L}\sigma^{z}_{i}$ to |$\ket{E}$ removes the $(-1)^{|f|}$ factor. Since the operator is a Clifford gate, the SRE-2 is unchanged. We immediately see that the SRE density of |$\ket{E}$ is given by half of the SRE density of |$\ket{\psi(\beta=0)}$, which is $m_{2}\approx 0.1733$. This result is within the errorbar of the numerical calculation in Ref. [87]. This highlights the wide applicability of the method employed here, as it is useful beyond states efficiently represented by low bond dimension tensor networks.
.2 Nonstabilizerness is always extensive
Figure S-1: (a) System-size dependence of $M_{2}$ near the Potts critical point, showing that the gauge-invariant nonstabilizerness is always an extensive quantity; (b) Power-law exponent $\alpha$ at different fitting ranges. Although a sublinear behavior is observed in the ordered phase, it flows towards a linear scaling in the asymptotic limit.
In this section, we demonstrate that $M_{2}$ is always an extensive quantity, even in ordered phases. In Fig. S-1(a), we can see that $M_{2}$ is always extensive with system-size $L$ near the Potts critical point and, therefore, the amount of non-Clifford resources needed to simulate the two distinguish phases always grows with $L$.
However, one can observe that $M_{2}$ has a sublinear dependence with $L$, as it is clear from Fig. 4(a) in the main text. This behavior is merely a finite-size effect, with the asymptotic scaling being a linear dependence with $L$. To illustrate this, we assume a power-law growth of $M_{2}$ ($M_{2}=AL^{\alpha}$) and show in Fig.S-1(b) how the power-law exponent $\alpha$ changes assuming different fitting ranges. In the ordered $\mathbb{Z}_{3}$ phase, it can be seen that the exponent grows as the system-size range of the fitting increases, and this suggests that the asymptotic behavior of $M_{2}$ always grows linearly with $L$.
The results in the other cuts analyzed throughout this paper are similar. Near the Tricritical Ising point, the SRE density $m_{2}$ becomes system size independent, showing that the gauge-invariant nonstabilizerness is a quantity that always grows linearly with system size, independent of the nature of the phase. The same conclusion can be obtained by looking at the third cut in the phase diagram where, despite some finite-size effects at small values of $L$, the curves of $m_{2}$ for $L=30$,$36$ and $42$ nearly overlap with each other.
.3 Sampling method and error analysis
The ground state $|\psi\rangle=\sum_{s}c_{s}|s\rangle$ of the U(1) QLM is obtained using exact diagonalization directly in its dual formulation (Eq. 3 of the main text). The Gauss law (nearest-neighbor constraint in the dual formulation) strongly reduces the growth of the Hilbert space dimension, which scales as $\mathrm{dim}\mathcal{H}_{L}=F_{L-1}+F_{L+1}$, where $F_{i}$ is the $i$-th Fibonacci number and $L$ is the number of matter sites. In the limit of large $L$, the Hilbert space dimension is given by $\mathrm{dim}\mathcal{H}_{L}\sim(1.618)^{L}$. Moreover, the periodic boundary conditions allow us to diagonalize the system directly within the momentum sector where the ground state resides. Therefore, by leveraging these two features, we can obtain the ground state of systems up to $L=42$ matter sites using Krylov space techniques.
Once we obtain the ground state in the momentum space, we transform it back to the real space. We then perform perfect sampling directly on its coefficients $|c_{s}|^{2}$ and compute all $\binom{N}{4}$ possibilities to choose from $N$ samples selected according to the cumulative distribution of the ground state coefficients (see Eq. 9 of the main text).
Although the gauge constraints eliminate many possibilities, we achieve a good numerical convergence by setting a sufficiently large number of samples and averaging over $50$ different realizations (see details below). The precision of our method is shown in Fig. 2(a) of the main text, where we benchmark the numerical results along the exact solvable line. It shows perfect agreement with the analytical prediction (black solid line), illustrating the robustness of our numerical algorithm. The convergence of $m_{2}$ with $N$ is relatively fast for moderate system sizes ($L=30$ qubits), as shown in Fig. 2(b), yielding errors of less than $1\%$ for any value of $N$ considered. Additionally, no finite-size effects are observed, with numerics matching the analytical prediction already for systems with $L=15$ spins (c.f. Fig. 2(c)).
Figure S-2: Boxplot of $50$ different runs along the exact solvable line ($V=2$) for $L=36$ qubits. Although different simulations yield different results, the difference between the mean value (red dashed line) and the exact solution (black dashed-dotted line) is less than $0.2\%$, showing the accuracy of our results.
It is important to stress that, because of the gauge constraint, a single Monte Carlo run may not be sufficient to obtain well-converged results. Therefore, to minimize the uncertainties in the computation of $m_{2}$, we perform $50$ independent runs and average the results. The statistical uncertainty is obtained by assuming that different runs are uncorrelated with each other, and is given by
$\sigma_{m_{2}}=\sqrt{\frac{\langle(m_{2}-\langle m_{2}\rangle)^{2}\rangle}{N_{\mathrm{real}}}}$ (S.2)
where $\langle m_{2}\rangle$ is the average over $N_{\mathrm{real}}$ different Monte Carlo runs. To benchmark the convergence of our numerical procedure, we set $N=500$ and evaluate the algorithm at a specific point along the solvable line (Eq. 7 of the main text) where the system has a large amount of nonstabilizerness ($m_{2}\approx 0.42$). The different results for $L=36$ are illustrated in Fig. S-2 in a boxplot representation, which conveniently shows how spread different runs can be. The difference between different runs is due to the Gauss law, which highly constrains the model so that the number of nonzero elements in the sum is much less than $\binom{N}{4}$ terms. However, the relative error between the mean value (red dashed lines) and the exact result (black dashed-dotted line) is approximately $0.16\%$, showing the robustness of our results.
It is important to mention that, throughout the letter, we consider at least $N=500$ samples. Therefore, despite the constraints imposed by the Gauss law, the statistical uncertainty is less than $1\%$, even for the largest system size available ($L=42$). However, our method loses its precision very quickly as $L$ grows, and one needs to consider $N>500$ samples for higher system sizes.